OpenStudy (abbycross167):

can someone please help me finish this question? I did it one but I have to revise it... I'll add my teachers response from the first submission. Question~ Describe the correlation between the U.S. shoe size and the European shoe size. My Answer~ The European size has a difference of 3 and 2. My teachers response~ Describe the correlation. How would I do that?

OpenStudy (abbycross167):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

So your teacher just repeated the question `Describe the correlation` and didn't say anything else?

OpenStudy (abbycross167):

yes sir that's all

OpenStudy (abbycross167):

ummm the question I had before this one may help you, if you would like me to show you it

jimthompson5910 (jim_thompson5910):

yes I'd like to see the full question and any other extra stuff needed (like a table of the data values)

OpenStudy (abbycross167):

ok 1 sec

OpenStudy (abbycross167):

Question~ 1. From the chart above, write the data as ordered pairs. (x,y) (1,31) (3,34) (5,39) (7,39) (9,41) (11,44) *question 2 was just a graph that I put the ordered pairs on* Question 3. Find an equation of the trend line (line of best fit) showing your work. Equation = y=3/2x+29.5 y1-y2 x1-x2 31-34 1 - 3 = -3=3 -2=2 M=3/2 31=3/2*1+b B=29.5 Equation = y=3/2x+29.5 Question 4. Describe the correlation between the U.S. shoe size and the European shoe size. The European size has a difference of 3 and 2.

OpenStudy (abbycross167):

@jim_thompson5910 did any of this help at all?

jimthompson5910 (jim_thompson5910):

ok let me look over it

jimthompson5910 (jim_thompson5910):

are you allowed to use a calculator? finding the line of best fit is a pain to do by hand

OpenStudy (abbycross167):

yes sir

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have?

OpenStudy (abbycross167):

A graphing calculator

jimthompson5910 (jim_thompson5910):

a texas instruments (TI) calculator?

OpenStudy (abbycross167):

this one~ Ti-84+ graphing calculator emulator ~ I got it from my school website

jimthompson5910 (jim_thompson5910):

ok that will work

jimthompson5910 (jim_thompson5910):

first enter the data given to you into the lists L1 and L2. Do you know how to do that?

OpenStudy (abbycross167):

no sir I don't, and I thought I already found the line of best fit?

jimthompson5910 (jim_thompson5910):

you only found the line through (1,31) and (3,34)

jimthompson5910 (jim_thompson5910):

hit the STAT key then go into the edit menu and hit enter

OpenStudy (abbycross167):

oh ok, but my teacher said that I got #3 right (I don't want to sound rude I just want to let you know what she said)

jimthompson5910 (jim_thompson5910):

ok then I guess some teachers are ok with using only 2 points, but just a heads up that isn't the true regression line anyways, let me re-think #4 then

jimthompson5910 (jim_thompson5910):

I guess one way is to graph the line and all of the points given to you. If that line is close to all of the points, then the correlation coefficient would be very close to +1. If not, then it will be closer to 0

OpenStudy (abbycross167):

oh ok, I'm really sorry if this is a trouble for you..

jimthompson5910 (jim_thompson5910):

no it's ok

OpenStudy (abbycross167):

ok so the answer to #4 would be "the correlation coefficient would be very close to +1. If not, then it will be closer to 0"

jimthompson5910 (jim_thompson5910):

well it depends on the graph. Did you graph the line and points on the same xy coordinate grid?

OpenStudy (abbycross167):

yes sir

jimthompson5910 (jim_thompson5910):

when I graphed it, I see that the line is fairly close to the other points. It definitely goes through (1,31) and (3,34) perfectly. It doesn't go through the other points, but it gets fairly close

jimthompson5910 (jim_thompson5910):

so this suggests the correlation coefficient is close to +1 meaning that the regression line is a fairly good fit. There is a strong correlation between the two sizes

OpenStudy (abbycross167):

oooh ok, thank you a lot!! you made this very easy!!

jimthompson5910 (jim_thompson5910):

no problem